1. A jet of water moving at a velocity of 30m/s impinges on a series of vanes moving at 15 m/s. The jet makes an angle 30o to the direction of motion of vanes at entry and leaves the vanes at 60o. Draw the inlet and outlet velocity diagrams and find i) vane angles at inlet and outlet ii) work done per second per Newton of water, and iii) hydraulic efficiency
1
Expert's answer
2021-05-12T08:38:28-0400

"v_1=30~\frac ms,~ v_2=15~\frac ms,~ \alpha=30°,~\beta=180°-60°=120°."

1.

"v_w=v_1\cos \alpha=25.98~\frac ms,"

"v_s=v_1\sin \alpha=15~\frac ms,"

"v_b=v_w-v_2=10.98~\frac ms,"

"v_r=\sqrt{v_b^2+v_s^2}=18.59~\frac ms,"

"\theta=\arctan \frac{v_s}{v_b}=53.8°,"

"\frac{v_r}{\sin(180°-\beta)}=\frac{v_2}{\sin(\beta-\psi)},"

"\psi=75.67°,"

"v_f=v_r\cos\psi-v_2=-10.4~\frac ms,"

2.

"W=\frac{(v_w+v_f)v_2}{g}=23.82~Nm/N ~weight~ of ~water,"

3.

"\eta=\frac{2(v_w+v_f)v_2}{v_1^2}=0.52=52\%."